A solution contains 0.375 M potassium hypochlorite and 0.435 M hypochlorous acid The pH of this solution is
Answer #1

=KClO is known as potassium hypochlorite whereas HOCl is Hypochlorous acid.

HOCl is weak acid and its conjugate base ClO- is strong.

KClO\rightarrow K^{+} + ClO^{-}

from the above equation, salt is a strong one that dissociates 1:1 that way potassium hypochlorite concentration = hypochlorite ion concentration

[ClO-] = 0.375

According to Henderson Hasselbalch equation, for which buffer solution contains weak acid and its conjugate base is

P^{H}= P^{Ka} + log([conjugate base]/[weak acid])

Ka = acid dissociation constant Pka = -log[Ka] which is for hypochlorous acid is 7.46 standard value

PH = 7.46 + log (0.375/0.435)

= 7.46 + log (0.862)

= 7.46+ log (8.62x10-1)

=7.46 +log(8.62) + log (10-1)

7.46+ 0.935-1

PH= 7.395